# An Elliott intertwining approach to classifying actions of $\mathrm{C}^{*}$-tensor categories

### 29 Jan 2024 v2

@@ 5.1. Intertwining through reparametrisation
Let $v_{n}=u_{1}u_{2}\ldots u_{n}$, we claim that $l^{X}(x):=\lim\limits_{n\to\infty}v_{n}\rhd h_{k_{n}}^{X}(x)\lhd v_{n}^{*}$ is well-defined for any $X\in\mathrm{Irr}(\mathcal{C})$ and any $x\in F(X)$. First note that $h^{X}$ is contractive by (iii) of Lemma 3.11, so $\limsup_{n\to\infty}\|h^{X}_{n}(x)\|\leq\|x\|$ for any $x\in F(X)$. In particular, the sequence $h^{X}_{k_{n}}$ is norm bounded.
We will now show that the sequence $(v_{n}\rhd h_{k_{n}}^{X}(x)\lhd v_{n}^{*})_{n\geq 1}$ is Cauchy for any $X\in\mathrm{Irr}(\mathcal{C})$ and any $x\in F(X)$. Given $\epsilon>0$, $n\in\mathbb{N}$, $X\in K_{n}$, and $x\in F(X)$, there exists $m\in\mathbb{N}$ and $y\in\mathcal{F}_{m}^{X}$ such that $\|x-y\|<\epsilon/3$. Then, for $j>l\geq m$, we may use (5.20) repeatedly, together with triangle inequality, to achieve that
$\|v_{j}\rhd h_{k_{j}}^{X}(y)\lhd v_{j}^{*}-v_{l}\rhd h_{k_{l}}^{X}(y)\lhd v_{l}^{*}\|$ $\leq\sum\limits_{i=l+1}^{j}2^{-i}<\epsilon/3$
for any $l$ greater than some large enough $N$. Moreover, since $h_{k_{l}}^{X}$ is a contractive linear map for any $l\in\mathbb{N}$ and any $X\in\mathrm{Irr}(\mathcal{C})$, it follows that $\|h_{k_{l}}^{X}(x)-h_{k_{l}}^{X}(y)\|<\epsilon/3$ Therefore,
$\|v_{j}\rhd h_{k_{j}}^{X}(x)\lhd v_{j}^{*}-v_{l}\rhd h_{k_{l}}^{X}(x)\lhd v_{l}^{*}\|<\epsilon,$
for all $l>N$, so the sequence is Cauchy. Thus, $l^{X}:F(X)\to G(X)$ is a well-defined map for every $X\in\mathrm{Irr}(\mathcal{C})$ and it is linear.
We let $\psi:A\to B$ be the linear map given by $\psi=l^{1_{\mathcal{C}}}$. If $v\in B_{\infty}$ is the unitary given by $(v_{n})_{n}$ and $\eta:\mathbb{N}\to\mathbb{N}$ is the map $\eta(n)=k_{n}$, we have that
$(\psi,l^{X})=(\mathrm{Ad}(v),h_{v}^{X})\circ(\eta^{*},r^{X})\circ(\phi,h^{X})$
for any $X\in\mathrm{Irr}(\mathcal{C})$. Hence, $\psi$ is a $*{}^{*}$-homomorphism and the pair $(\psi,l)$ induces a cocycle morphism. Extending by linearity, we can define $l^{X}$ in the same way for any $X\in\mathcal{C}$ and it is straightforward to see that these maps are linear. Moreover, $(\psi,l)$ is approximately unitarily equivalent to $(\phi,h)$, which finishes the proof. ∎
@@
As noted in Remark 2.7 the assumption that $G(X)$ are algebraically finitely generated in Theorem 5.9 is immediate whenever the acting category $\mathcal{C}$ is a unitary tensor category.
@@ 6. One sided intertwining arguments
We start this section by showing a tensor category equivariant adaptation of the classical one-sided intertwining argument (see [38, Proposition 2.3.5]). A group equivariant version of this result can be found in [40, Proposition 4.3]. We end this section by proving an asymptotic Elliott two-sided intertwining (see Theorem 6.4).
@@
Let $\mathcal{C}$ be a semisimple $*{}^{*}$-tensor category with countably many isomorphism classes of simple objects. Let $(F,J):\mathcal{C}\curvearrowright A$ and $(G,I):\mathcal{C}\curvearrowright B$ be actions on separable $*{}^{*}$-algebras and $(\varphi,h):(A,F,J)\rightarrow(B,G,I)$ an injective cocycle morphism. Then $(\varphi,h)$ is asymptotically unitarily equivalent to a cocycle conjugacy if and only if:
@@
    $\mathrm{dist}(z_{1}^{*}\rhd y\lhd z_{1},h^{X}(F(X)))\leq\varepsilon$ for all $X\in K$ and $y\in\mathcal{G}^{X}$.
@@
The proof of this Theorem will follow the proof of [40, Proposition 4.3] closely. We start by showing the “only if” statement. Let $(\Phi,H)$ be a cocycle conjugacy such that $(\Phi,H)\cong_{u}(\varphi,h)$. By Lemma 4.18 there exists a strictly continuous map $w:[0,\infty)\rightarrow\mathcal{U}(\mathcal{M}(B))$ such that
@@
We claim that the unitary path $z_{t}=w_{n_{1}}^{*}w_{(1-t)n_{1}+tn_{2}}$ satisfies A.1 and A.2. Indeed, it is a strictly continuous path $z:[0,1]\rightarrow\mathcal{U}(\mathcal{M}(B))$ with $z_{0}=1$. Moreover, for $X\in K$, $t\in[0,1]$ and $x\in\mathcal{F}^{X}$, using (6.2) we have that
$\|h^{X}(x)-z_{t}$ $*\displaystyle\rhd h^{X}(x)\lhd z_{t}^{*}\|$
@@
This shows condition A.1. We now turn to condition A.2. Let $X\in K$, $y\in\mathcal{G}^{X}$, and $x=(H^{X})^{-1}(w_{n_{1}}\rhd y\lhd w_{n_{1}}^{*})$. Then, we get that
@@
(6.3)ϵ.superscriptitalic-(6.3italic-)absentitalic-${{\leq}}\epsilon.$
As $x\in F(X)$, condition A.2 holds.
We now turn to the if direction. To prove this statement we will use A.1-A.2 to construct a strictly continuous path of unitaries $v_{t}\in\mathcal{U}(\mathcal{M}(B))$ for $t\in[0,\infty)$ such that $(\mathrm{Ad}(v_{t}),h_{v_{t}})\circ(\varphi,h)$ converges to a cocycle conjugacy.
Let $\{x_{n}^{X}\}_{n\in\mathbb{N}}\subset F(X)$, $\{y_{n}^{X}\}_{n\in\mathbb{N}}\subset G(X)$ be countable dense subsets for any $X\in\mathrm{Irr}(\mathcal{C})$ and $K_{n}$ increasing finite subsets of $\mathrm{Irr}(\mathcal{C})$ containing $1_{\mathcal{C}}$ such that $\mathrm{Irr}(\mathcal{C})=\bigcup_{n\in\mathbb{N}}K_{n}$. Firstly, use A.1-A.2 to find $x_{1,1}^{X}\in F(X)$ for $X\in K_{1}$ and $z^{(1)}:[0,1]\rightarrow\mathcal{U}(\mathcal{M}(B))$ such that $z_{0}^{(1)}=1$ and for $0\leq t\leq 1$
@@
Again use A.1-A.2 to find $x_{2,1}^{X}$, $x_{2,2}^{X}$ in $F(X)$ for $X\in K_{2}$, $z^{(2)}:[0,1]\rightarrow\mathcal{U}(\mathcal{M}(B))$ such that $z^{(2)}_{0}=1$ and for every $0\leq t\leq 1$
@@
    $\|(z_{1}^{(n)})^{*}\ldots(z_{1}^{(1)})^{*}\rhd y_{j}^{X}\lhd z_{1}^{(1)}\ldots z_{1}^{(n)}-h^{X}(x_{n,j}^{X})\|\leq 2^{-n}$ for $X\in K_{n}$ and $1\leq j\leq n$.
Then use A.1-A.2 to get $\{x_{n+1,j}^{X}\}_{j\leq n+1}\in F(X)$ for $X\in K_{n+1}$ and $z^{(n+1)}:[0,1]\rightarrow\mathcal{U}(\mathcal{M}(B))$ such that for all $t\in[0,1]$
@@
    $\|(z_{1}^{(n+1)})^{*}\ldots(z_{1}^{(1)})^{*}\rhd y_{j}^{X}\lhd z_{1}^{(1)}\ldots z_{1}^{(n+1)}-h^{X}(x_{n+1,j}^{X})\|\leq 2^{-(n+1)}$ for $X\in K_{n+1}$ and $11\leq j\leq n+$ 1.
We carry on inductively to construct $z^{(n)}$ and $x_{m,j}$ for every $n,m\in\mathbb{N}$ and $j\leq m$ satisfying S.1-S.3. We may now define the path $v_{t}:[0,\infty)\rightarrow\mathcal{U}(\mathcal{M}(B))$ by $v_{t}=z_{1}^{(1)}\ldots z_{1}^{(n)}z_{t-n}^{(n+1)}$ for every $t\in[n,n+1]$. This path is strictly continuous on every open interval $(n,n+1)$ for $n\in\mathbb{N}$ as the paths $z^{(k)}$ are strictly continuous for each $k\in\mathbb{N}$. Moreover, the path $v_{t}$ is strictly continuous at each $n\in\mathbb{N}$ as $z_{0}^{(k)}=1$ for every $k\in\mathbb{N}$. Adjoining by $v_{t}$ we obtain a continuous path of cocycle morphisms $(\mathrm{Ad}(v_{t}),h_{v_{t}})\circ(\varphi,h)=(\psi_{t},h_{t})$ where $\psi_{t}=\mathrm{Ad}(v_{t})\circ\varphi$ and $h_{t}^{X}=h_{v_{t}}^{X}\circ h^{X}$ for any $t\in[0,\infty)$ and $X\in\mathrm{Irr}(\mathcal{C})$.
We claim that as $t\rightarrow\infty$ the path $\psi_{t}$ converges to an isomorphism $\Psi$ and that the path $h_{t}^{X}$ converges to a bijective linear map $H^{X}$ for all $X\in\mathrm{Irr}(\mathcal{C})$ such that the pair $(\Psi,H)$ induces a cocycle morphism (recall Remark 3.12). For any $X\in\mathrm{Irr}(\mathcal{C})$ and $j\in\mathbb{N}$ the net $(h_{t}^{X}(x_{j}^{X}))_{t\geq 0}$ is Cauchy by S.1. Since the set $\{x_{j}^{X}\}_{j\in\mathbb{N}}$ is dense in $F(X)$ for any $X\in\mathrm{Irr}(\mathcal{C})$ a standard triangle inequality argument shows that the net $(h_{t}^{X}(x))_{t\geq 0}$ converges for any $x\in F(X)$ and $X\in\mathrm{Irr}(\mathcal{C})$. We may hence define $H^{X}(x)=\lim_{t\rightarrow\infty}h_{t}^{X}(x)$ for all $x\in F(X)$ and $X\in\mathrm{Irr}(\mathcal{C})$. The maps $H^{X}:F(X)\rightarrow G(X)$ are linear as inherited by the linearity of $h_{t}^{X}$. Letting $X=1_{\mathcal{C}}$ the map $h^{1_{\mathcal{C}}}$ coincides with $\varphi$ and we get that $\Psi(a):=\lim_{t\rightarrow\infty}\mathrm{Ad}(v_{t})\circ\varphi(a)=\lim_{t\rightarrow\infty}\psi_{t}(a)$ exists for all $a\in A$. As each $\psi_{t}$ is a $*{}^{*}$-homomorphism, then so is $\Psi$. In light of Remark 3.12, it suffices to check that the family of maps $\{H^{X}\}_{X\in\mathrm{Irr}(\mathcal{C})}$ satisfies the conditions (i)-(v) of Lemma 3.11 to conclude that $(\Psi,H)$ is a cocycle morphism. Condition (v) follows by construction, whereas conditions (i)-(iv) are easily verified as $(\psi_{t},h_{t})$ is a cocycle morphism for every $t\in[0,\infty)$ and one may approximate $\Psi$ pointwise by $\psi_{t}$ and $H$ pointwise by $h_{t}$ for some large enough $t$.
It remains to check that $\Phi$ is bijective and $H^{X}$ is bijective for every $X\in\mathcal{C}$. By Remark 3.12 and by taking $X=1_{\mathcal{C}}$, it suffices to check that $H^{X}$ is bijective for every $X\in\mathrm{Irr}(\mathcal{C})$. Firstly, each $H^{X}$ is isometric. To see this, note that $\Psi$ is an injective $*{}^{*}$-homomorphism and hence isometric, so as $(\Psi,H)$ is a cocycle morphism
@@
$\leq 2^{1-n},$
for any $n\geq n_{0}$. So $\|y_{j}^{X}-H^{X}(x_{n,j}^{X})\|\leq 2^{1-n}+2^{-n}$. Therefore, as $n$ may be chosen arbitrarily and $\{y_{j}^{X}\}_{j\in\mathbb{N}}$ is dense in $G(X)$, it follows that $H^{X}$ is surjective. ∎
@@
Let $\mathcal{C}$ be a semisimple $*{}^{*}$-tensor category with countably many isomorphism classes of simple objects. Let $(F,J):\mathcal{C}\curvearrowright A$ and $(G,I):\mathcal{C}\curvearrowright B$ be actions on separable $*{}^{*}$-algebras and $(\varphi,h):(A,F,J)\rightarrow(B,G,I)$ an injective cocycle morphism. Then $(\varphi,h)$ is approximately unitarily equivalent to a cocycle conjugacy if and only if:
@@
Let $\mathcal{C}$ be a semisimple $*{}^{*}$-tensor category with countably many isomorphism classes of simple objects. Let $(F_{n},J^{(n)}):\mathcal{C}\curvearrowright A_{n}$ be a sequence of actions on separable $\mathrm{C}^{*}$-algebras. Let
@@
    $\mathrm{dist}(y_{1}^{*}\rhd\eta\lhd y_{1},h_{1,n}^{X}(F_{1}(X)))\leq\varepsilon$ for any $X\in K$, $\eta\in\mathcal{F}_{n}^{X}$
@@
Let $\mathcal{C}$ be a semisimple $*{}^{*}$-tensor category with countably many isomorphism classes of simple objects. Let $(F,J):\mathcal{C}\curvearrowright A$ and $(G,I):\mathcal{C}\curvearrowright B$ be actions on separable $\mathrm{C}^{*}$-algebras. Let
@@
[…]
@@
Furthermore, the cocycle morphism $(\kappa,r)$ is asymptotically unitarily equivalent to the cocycle morphism induced by the identity map on $A$, so Lemma 6.3 gives that $(\kappa_{\infty},r_{\infty})$ is asymptotically unitarily equivalent to a cocycle conjugacy $(K,R):(A,F,J)\to(A^{(\infty)},F^{(\infty)},J^{(\infty)})$. Likewise, $(\theta_{\infty},s_{\infty})$ is asymptotically unitarily equivalent to a cocycle conjugacy $(\Theta,S):(B,G,I)\to(B^{(\infty)},G^{(\infty)},I^{(\infty)})$.

### 08 Apr 2026 v3

@@ 6.1. Intertwining through reparametrisation
Let $v_{n}=u_{1}u_{2}\ldots u_{n}$, we claim that $l^{X}(x):=\lim\limits_{n\to\infty}v_{n}\rhd h_{k_{n}}^{X}(x)\lhd v_{n}^{*}$ is well-defined for any $X\in\mathrm{Irr}(\mathcal{C})$ and any $x\in F(X)$. First let $y\in\bigcup_{n\in\mathbb{N}}\mathcal{F}_{n}^{X}$ for $X\in\mathrm{Irr}(\mathcal{C})$ and $\varepsilon>0$. Then there exists $n\in\mathbb{N}$ such that $X\in K_{n}$, $y\in\mathcal{F}_{n}^{X}$ and $2^{-n}<\varepsilon$. In particular, for any $j>l\geq n$ we may use (6.20) repeatedly, together with triangle inequality, to achieve that
$\|v_{j}\rhd h_{k_{j}}^{X}(y)\lhd v_{j}^{*}-v_{l}\rhd h_{k_{l}}^{X}(y)\lhd v_{l}^{*}\|$ $\leq\sum\limits_{i=l+1}^{j}2^{-i}<\epsilon.$
As $\varepsilon$ is arbitrary $v_{n}\rhd h_{k_{n}}^{X}(y)\lhd v_{n}^{*}$ is Cauchy for any $X\in\mathrm{Irr}(\mathcal{C})$ and any $y\in\bigcup_{n\in\mathbb{N}}\mathcal{F}_{n}^{X}$. This implies that, letting $(v_{n})_{n}=V\in\mathcal{U}(B_{\infty})$, $\eta:\mathbb{N}\rightarrow\mathbb{N}$ be given by $\eta(n)=k_{n}$ and their induced cocycle morphisms by $(\mathrm{Ad}(V),h_{V})$ and $(\eta^{*},r)$,
$(\mathrm{Ad}(V),h_{V})\circ(\eta^{*},r)\circ(\phi,h):(A,F,J)\rightarrow(B_{\infty},G_{\infty},I^{\infty})$
is a cocycle morphism such that it sends a dense subspace of $F(X)$ into the closed subspace $G(X)\subset G(X)_{\infty}$ for each $X\in\mathrm{Irr}(\mathcal{C})$. Thus
$(h_{V})^{X}\circ r^{X}\circ h^{X}:F(X)\to G(X)\subset G(X)_{\infty}$
by continuity for any $X\in\mathrm{Irr}(\mathcal{C})$. This precisely shows that
$\lim_{n\rightarrow\infty}v_{n}\rhd h_{k_{n}}^{X}(x)\lhd v_{n}^{*}$
exists for all $x\in F(X)$. Let $\psi:A\to B$ be the linear map given by $\psi=l^{1_{\mathcal{C}}}$. By construction we have that
$(\psi,l)=(\mathrm{Ad}(V),h_{V})\circ(\eta^{*},r)\circ(\phi,h)$
for any $X\in\mathrm{Irr}(\mathcal{C})$. Hence, $\psi$ is a -homomorphism and the pair $(\psi,l)$ induces a cocycle morphism. Extending by linearity, we can define $l^{X}$ in the same way for any $X\in\mathcal{C}$ and it is straightforward to see that these maps are linear. Moreover, $(\psi,l)$ is approximately unitarily equivalent to $(\phi,h)$. Finally, to show that the approximate unitary equivalence of $(\psi,l)$ and $(\phi,h)$ implies unitary equivalence follows as in [18, Lemma 4.1]. ∎
@@
Note that in Theorem 6.8, if the approximate unitary equivalences are realised by unitaries in the minimal unitisations, then we may drop the assumption of unitality on $B$. Moreover, as noted in Remark 2.9, the assumption that $G(X)$ are finite rank in Theorem 6.8 is immediate whenever the acting category $\mathcal{C}$ is a unitary tensor category and $B$ is unital.
@@ 7. One sided intertwining arguments
We start this section by showing a tensor category equivariant adaptation of the classical one-sided intertwining argument (see [40, Proposition 2.3.5]). A group equivariant version of this result can be found in [41, Proposition 4.3]. We end this section by proving an asymptotic Elliott two-sided intertwining (see Theorem 7.4). First note that as $\mathrm{Ad}(v)$ is an isometric -homomorphism for any unitary $v\in\mathcal{U}(\mathcal{M}(A))$, injectivity of a cocycle morphism, which coincides with the injectivity of it’s underlying -homomorphism, is preserved under both approximate and asymptotic unitary equivalence.
@@
Let $\mathcal{C}$ be a semisimple C-tensor category with countably many isomorphism classes of simple objects. Let $(F,J):\mathcal{C}\curvearrowright A$ and $(G,I):\mathcal{C}\curvearrowright B$ be actions on separable C-algebras and $(\varphi,h):(A,F,J)\rightarrow(B,G,I)$ an injective cocycle morphism. Then $(\varphi,h)$ is asymptotically unitarily equivalent to a cocycle conjugacy if and only if:
@@
    $\mathrm{dist}(z_{1}^{*}\rhd y\lhd z_{1},h^{X}(F(X)))\leq\varepsilon$ for all $X\in K$ and $y\in\mathcal{G}^{X}$. [^18] Note that the injectivity assumption is redundant for the only if statement as it follows from the hypothesis. However, this assumption it is still required for the if direction.
@@
The proof of this Theorem will follow the proof of [41, Proposition 4.3] closely. We start by showing the “only if” statement. Let $(\Phi,H)$ be a cocycle conjugacy such that $(\Phi,H)\cong_{u}(\varphi,h)$. By Lemma 5.12 there exists a strictly continuous map $w:[0,\infty)\rightarrow\mathcal{U}(\mathcal{M}(B))$ such that
@@
We claim that the unitary path $z_{t}=w_{n_{1}}^{*}w_{(1-t)n_{1}+tn_{2}}$ satisfies A.1 and 18. Indeed, it is a strictly continuous path $z:[0,1]\rightarrow\mathcal{U}(\mathcal{M}(B))$ with $z_{0}=1$. Moreover, for $X\in K$, $t\in[0,1]$ and $x\in\mathcal{F}^{X}$, using (7.2) we have that
$\|h^{X}(x)-z_{t}$ $\rhd h^{X}(x)\lhd z_{t}^{*}\|$
@@
This shows condition A.1. We now turn to condition 18. Let $X\in K$, $y\in\mathcal{G}^{X}$, and $x=(H^{X})^{-1}(w_{n_{1}}\rhd y\lhd w_{n_{1}}^{*})$. Then, we get that
@@
[…]
As $x\in F(X)$, condition 18 holds.
We now turn to the if direction. To prove this statement we will use A.1-18 to construct a strictly continuous path of unitaries $v_{t}\in\mathcal{U}(\mathcal{M}(B))$ for $t\in[0,\infty)$ such that $(\mathrm{Ad}(v_{t}),h_{v_{t}})\circ(\varphi,h)$ converges to a cocycle conjugacy.
Let $\{x_{n}^{X}\}_{n\in\mathbb{N}}\subset F(X)$, $\{y_{n}^{X}\}_{n\in\mathbb{N}}\subset G(X)$ be countable dense subsets for any $X\in\mathrm{Irr}(\mathcal{C})$ and $K_{n}$ increasing finite subsets of $\mathrm{Irr}(\mathcal{C})$ containing $1_{\mathcal{C}}$ such that $\mathrm{Irr}(\mathcal{C})=\bigcup_{n\in\mathbb{N}}K_{n}$. Firstly, use A.1-18 to find $x_{1,1}^{X}\in F(X)$ for $X\in K_{1}$ and $z^{(1)}:[0,1]\rightarrow\mathcal{U}(\mathcal{M}(B))$ such that $z_{0}^{(1)}=1$ and for $0\leq t\leq 1$
@@
Again use A.1-18 to find $x_{2,1}^{X}$, $x_{2,2}^{X}$ in $F(X)$ for $X\in K_{2}$, $z^{(2)}:[0,1]\rightarrow\mathcal{U}(\mathcal{M}(B))$ such that $z^{(2)}_{0}=1$ and for every $0\leq t\leq 1$
@@
    $\lVert\left((z_{1}^{(n)})^{*}\ldots(z_{1}^{(1)})^{*}\rhd y_{j}^{X}\lhd z_{1}^{(1)}\ldots z_{1}^{(n)}\right)-h^{X}(x_{n,j}^{X})\rVert\leq 2^{-n}$ for $X\in K_{n}$ and $1\leq j\leq n$.
Then use A.1-18 to get $\{x_{n+1,j}^{X}\}_{j\leq n+1}\in F(X)$ for $X\in K_{n+1}$ and $z^{(n+1)}:[0,1]\rightarrow\mathcal{U}(\mathcal{M}(B))$ such that for all $t\in[0,1]$
@@
    $\|\left((z_{1}^{(n+1)})^{*}\ldots(z_{1}^{(1)})^{*}\rhd y_{j}^{X}\lhd z_{1}^{(1)}\ldots z_{1}^{(n+1)}\right)-h^{X}(x_{n+1,j}^{X})\|\leq 2^{-(n+1)}$ for $X\in K_{n+1}$ and $1\leq j\leq n+1$.
We carry on inductively to construct $z^{(n)}$ and $x_{m,j}^{X}$ for every $n,m\in\mathbb{N}$ and $j\leq m$ satisfying S.1-S.3. We may now define the path $v_{t}:[0,\infty)\rightarrow\mathcal{U}(\mathcal{M}(B))$ by $v_{t}=z_{1}^{(1)}\ldots z_{1}^{(n)}z_{t-n}^{(n+1)}$ for every $t\in[n,n+1]$. This path is strictly continuous on every open interval $(n,n+1)$ for $n\in\mathbb{N}$ as the paths $z^{(k)}$ are strictly continuous for each $k\in\mathbb{N}$. Moreover, the path $v_{t}$ is strictly continuous at each $n\in\mathbb{N}$ as $z_{0}^{(k)}=1$ for every $k\in\mathbb{N}$. Adjoining by $v_{t}$ we obtain a continuous path of cocycle morphisms $(\mathrm{Ad}(v_{t}),h_{v_{t}})\circ(\varphi,h)=(\psi_{t},h_{t})$ where $\psi_{t}=\mathrm{Ad}(v_{t})\circ\varphi$ and $h_{t}^{X}=h_{v_{t}}^{X}\circ h^{X}$ for any $t\in[0,\infty)$ and $X\in\mathrm{Irr}(\mathcal{C})$.
We claim that as $t\rightarrow\infty$ the path $\psi_{t}$ converges to an isomorphism $\Psi$ and that the path $h_{t}^{X}$ converges to a bijective linear map $H^{X}$ for all $X\in\mathrm{Irr}(\mathcal{C})$ such that the pair $(\Psi,H)$ induces a cocycle morphism (recall Remark 3.12). For any $X\in\mathrm{Irr}(\mathcal{C})$ and $j\in\mathbb{N}$ the net $(h_{t}^{X}(x_{j}^{X}))_{t\geq 0}$ is Cauchy by S.1. Since the set $\{x_{j}^{X}\}_{j\in\mathbb{N}}$ is dense in $F(X)$ for any $X\in\mathrm{Irr}(\mathcal{C})$ a standard triangle inequality argument shows that the net $(h_{t}^{X}(x))_{t\geq 0}$ converges for any $x\in F(X)$ and $X\in\mathrm{Irr}(\mathcal{C})$. We may hence define $H^{X}(x)=\lim_{t\rightarrow\infty}h_{t}^{X}(x)$ for all $x\in F(X)$ and $X\in\mathrm{Irr}(\mathcal{C})$. The maps $H^{X}:F(X)\rightarrow G(X)$ are linear as inherited by the linearity of $h_{t}^{X}$. In light of Remark 3.12, it suffices to check that the family of maps $\{H^{X}\}_{X\in\mathrm{Irr}(\mathcal{C})}$ satisfies the conditions (ii)-(iv) of Lemma 3.10 to conclude that $(\Psi,H)$ is a cocycle morphism. They are easily verified as $(\psi_{t},h_{t})$ is a cocycle morphism for every $t\in[0,\infty)$ and one may approximate $H$ pointwise by $h_{t}$ for some large enough $t$.
It remains to check that $H^{X}$ is bijective for every $X\in\mathcal{C}$. By Remark 3.12, it suffices to check that $H^{X}$ is bijective for every $X\in\mathrm{Irr}(\mathcal{C})$. Firstly, each $H^{X}$ is isometric. To see this, note that $\Psi$ is an injective -homomorphism and hence isometric, so as $(\Psi,H)$ is a cocycle morphism
@@
$=2^{-n},$
for any $n\geq n_{0}$. So $\|y_{j}^{X}-H^{X}(x_{n,j}^{X})\|\leq 2^{1-n}$. Therefore, as $n$ may be chosen arbitrarily and $\{y_{j}^{X}\}_{j\in\mathbb{N}}$ is dense in $G(X)$, it follows that $H^{X}$ is surjective. ∎
@@
Let $\mathcal{C}$ be a semisimple C-tensor category with countably many isomorphism classes of simple objects. Let $(F,J):\mathcal{C}\curvearrowright A$ and $(G,I):\mathcal{C}\curvearrowright B$ be actions on separable C-algebras and $(\varphi,h):(A,F,J)\rightarrow(B,G,I)$ an injective cocycle morphism. Then $(\varphi,h)$ is approximately unitarily equivalent to a cocycle conjugacy if and only if:
@@
Let $\mathcal{C}$ be a semisimple C-tensor category with countably many isomorphism classes of simple objects. Let $(F_{n},J^{(n)}):\mathcal{C}\curvearrowright A_{n}$ be a sequence of actions on separable $\mathrm{C}^{*}$-algebras. Let
@@
    $\mathrm{dist}(y_{1}^{*}\rhd\eta\lhd y_{1},h_{1,n}^{X}(F_{1}(X)))\leq\varepsilon$ for any $X\in K$, $\eta\in\mathcal{F}_{n}^{X}$ .
@@
Let $\mathcal{C}$ be a semisimple C-tensor category with countably many isomorphism classes of simple objects. Let $(F,J):\mathcal{C}\curvearrowright A$ and $(G,I):\mathcal{C}\curvearrowright B$ be actions on separable $\mathrm{C}^{*}$-algebras. Let
@@
(7.4) $\dots\ignorespaces\ignorespaces\ignorespaces\ignorespaces$${F(X)\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{h^{X}}$$\scriptstyle{r^{X}}$${F(X)\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{h^{X}}\textstyle{\dots}\textstyle{\dots\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${G(X)\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{l^{X}}$$\scriptstyle{s^{X}}$${G(X)\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{l^{X}}$$\scriptstyle{s^{X}}$${\dots\quad.}$
@@
Furthermore, the cocycle morphism $(\kappa,r)$ is asymptotically unitarily equivalent to the cocycle morphism induced by the identity map on AA, which in particular implies that $(\kappa,r)$ is injective and thus Lemma 7.3 gives that $(\kappa_{\infty},r_{\infty})$ is asymptotically unitarily equivalent to a cocycle conjugacy $(K,R):(A,F,J)\to(A^{(\infty)},F^{(\infty)},J^{(\infty)})$. Likewise, $(\theta_{\infty},s_{\infty})$ is asymptotically unitarily equivalent to a cocycle conjugacy $(\Theta,S):(B,G,I)\to(B^{(\infty)},G^{(\infty)},I^{(\infty)})$.
@@
Remark 7.5.
Note that in Theorem 7.4, if the asymptotic unitary equivalences are realised by unitaries in the minimal unitisations, then we may drop the assumption of extendibility on the connecting morphisms. This follows as extendibility is only required to evaluate a morphism on a given unitary in the multiplier algebra. However, there is no issue in doing so when the unitary is in the minimal unitisation.